# On Discounted Infinite-Time Mean Field Games

### v1

@@ Abstract
We construct a Nash equilibrium using the stochastic maximum principle and infinite-time forward-backward stochastic differential equations. By employing the elliptic master equation, we provide a representation for the Nash equilibrium.
Keywords. discounted infinite-time mean field games, infinite-time forward-backward equations, elliptic master equation.
@@ 1 Introduction
The investigation of general nonlinear BSDEs was pioneered by Pardoux and Peng [12, 13] in the early 1990s, which is now a typical tool in stochastic optimization problems. We shall solve the mean field game problem using the Pontryagin’s maximum principle and infinite-time forward-backward stochastic differential equations. The foundational work in [14] proved the existence and uniqueness of solutions to the infinite-time FBSDEs, later extended by [1] to the Mckean-Vlasov FBSDEs, which play crucial roles in solving equilibrium solutions for mean field games. In this paper, we partition all agents into a representative player and the social equilibrium, and characterize the equilibrium state through the following system of infinite-time FBSDE:
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\xi+b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (1.5)
$\begin{cases}\mathrm{d}X_{t}^{x}=x+b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (1.6)
@@ 2 Preliminaries
$W_{p}(\mu,\nu)\triangleq\inf\Big{\{}\big{(}\mathbb{E}[|\xi-\eta|^{q}]\big{)}^{1/q}:~{}\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big{\}}.$
@@ 3 Infinite-time mean field games with discounting
We observe that the cost function depends exclusively on the distribution of $X_{t}^{\xi,\alpha}$ and we have
$\xi^{\prime}\in\mathbb{L}^{2}(\mathcal{F}_{0}),\quad\mathcal{L}_{\xi^{\prime}}=\mathcal{L}_{\xi}\quad\Longrightarrow\quad J(x,\xi^{\prime};\alpha,\beta)=J(x,\xi;\alpha,\beta),\quad~{}\mbox{for all $x,\alpha,\beta$.}$ (3.9)
Therefore, we may define
$J(x,\mu;\alpha,\beta)\triangleq J(x,\xi;\alpha,\beta),\quad\quad\xi\in\mathbb{L}^{2}(\mathcal{F}_{0},\mu).$ (3.10)
Now, for any $(x,\mu)\in\mathbb{R}\times{\cal P}_{2}$ and $\alpha\in\mathcal{A}$, we consider the infimum
$V(x,\mu;\alpha)\triangleq\inf_{\beta\in\mathcal{A}}J(x,\mu;\alpha,\beta).$ (3.11)
@@
We say a function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ is a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if the closed-loop controls $\alpha_{s}^{\mu_{0}}={\alpha}^{*}(X_{s}^{\xi,\alpha^{\mu_{0}}},\mathcal{L}_{X_{s}^{\xi,\alpha^{\mu_{0}}}})$,$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi,\alpha^{\mu_{0}}}})$ satisfy
$\alpha^{x}\in\arg\min_{\beta\in\mathcal{A}}J(x,\mu_{0};\alpha^{\mu_{0}},\beta).$ (3.12)
When this Nash equilibrium $\alpha^{*}$ exists, we can define the value function with respect to this Nash equilibrium as:
$V(x,\mu_{0})\triangleq J(x,\mu_{0};\alpha^{\mu_{0}},\alpha^{x}).$ (3.13)
@@ 4.1 Pontryagin’s maximum principle
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\xi+b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (4.29)
$\begin{cases}\mathrm{d}X_{t}^{x}=x+b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (4.30)
@@ 4.2 Solvability of mean field game FBSDEs
If we set $f_{1}(x,a)=Ax^{2}+Ca^{2}$, we have $\iota=A,\eta=C,\zeta=2C$. Then the requirement in (4.43) becomes
@@ 5 Master equation representation
Assume the master equation (1.7) admits a bounded classical solution $U(x,\mu)\in{\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ with bounded derivatives, and the following SDEs admit unique solutions in $L^{2}_{r}$ for each $x\in\mathbb{R}$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu)$:
@@
Since $U$ and its derivatives are bounded, taking the limit as $T\rightarrow+\infty$ yields

### v2

@@ Abstract
We construct a Nash equilibrium using the stochastic maximum principle and infinite-time forward-backward stochastic differential equations. By employing the elliptic master equations, a class of distribution-dependent elliptic PDEs , we provide a representation for the Nash equilibrium. And we prove that the solutions to a system of infinite-time forward-backward stochastic differential equations can be employed to construct viscosity solutions for a class of distribution-dependent elliptic PDEs.
Keywords. discounted infinite-time mean field games, infinite-time forward-backward equations, elliptic master equations.
@@ 1 Introduction
The investigation of general nonlinear BSDEs was pioneered by Pardoux and Peng [12, 13] in the early 1990s, which is now a typical tool in stochastic optimization problems. We shall solve the mean field game problem using the Pontryagin’s maximum principle and infinite-time forward-backward stochastic differential equations. The foundational work in [14] proved the existence and uniqueness of solutions to the infinite-time FBSDEs, and subsequent work in [15] investigated a more general setting and established connections with quasilinear elliptic PDEs. Recently, [1] extended this framework to the Mckean-Vlasov FBSDEs, which play crucial roles in solving equilibrium solutions for mean field games. In this paper, we partition all agents into a representative player and the social equilibrium, and characterize the equilibrium state through the following system of infinite-time FBSDE:
$\begin{cases}\mathrm{d}X_{t}^{\xi}=b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (1.5)
$\begin{cases}\mathrm{d}X_{t}^{x}=b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (1.6)
@@
This paper is organized as follows: in section 2, we present the preliminaries of problems in this paper; in section 3, we introduce the infinite-time mean field games with discounting and the definition of Nash equilibrium; in section 4 we characterize the equilibrium states through a system of infinite-time FBSDEs and in section 5 we introduce the elliptic master equation to provide a representation for the Nash equilibrium; in section 6, we provide a viscosity solution for distribution-dependent elliptic PDEs by virtue of the class of FBSDEs introduced in section 4.
@@ 2 Preliminaries
$W_{p}(\mu,\nu):=\inf\Big{\{}\big{(}\mathbb{E}[|\xi-\eta|^{q}]\big{)}^{1/q}:~{}\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big{\}}.$
@@ 3 Infinite-time mean field games with discounting
For any $(x,\xi)\in\mathbb{R}\times\mathbb{L}^{2}(\mathcal{F}_{0})$ and $\alpha\in\mathcal{A}$, we consider the infimum
$V(x,\xi;\alpha)\triangleq\inf_{\beta\in\mathcal{A}}J(x,\xi;\alpha,\beta).$ (3.9)
@@
We say a Lipschitz function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ is a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if for any initial state $\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with distribution $\mu_{0}$, the closed-loop controls $\alpha_{s}^{\xi_{0}}={\alpha}^{*}(X_{s}^{\xi_{0},\alpha^{\xi_{0}}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$,$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$ satisfy
$\alpha^{x}\in\arg\min_{\beta\in\mathcal{A}}J(x,\xi_{0};\alpha^{\xi_{0}},\beta).$ (3.10)
When this Nash equilibrium $\alpha^{*}$ exists, we have
$\xi,\xi^{\prime}\in\mathbb{L}^{2}(\mathcal{F}_{0}),\quad\mathcal{L}_{\xi^{\prime}}=\mathcal{L}_{\xi}\quad\Longrightarrow\quad\mathcal{L}_{X_{t}^{\xi,\alpha^{\xi}}}=\mathcal{L}_{X_{t}^{\xi^{\prime},\alpha^{\xi^{\prime}}}},\quad~{}\mbox{for a.e. $t\geq 0$.}$ (3.11)
Therefore, we can define
$V(x,\mu)\triangleq J(x,\xi_{0};\alpha^{\xi_{0}},\alpha^{x}),\quad\quad\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0},\mu).$ (3.12)
@@ 4.1 Pontryagin’s maximum principle
$\begin{cases}\mathrm{d}X_{t}^{\xi}=b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (4.12)
$\begin{cases}\mathrm{d}X_{t}^{x}=b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (4.13)
@@ 4.2 Solvability of mean field game FBSDEs
If we set $f_{1}(x,a)=Ax^{2}+Ca^{2},A>0,C>0$, we have $\iota=A,\eta=C,\zeta=2C$. Then the requirement in (4.26) becomes
@@ 5 Master equation representation
Assume the master equation (1.7) admits a classical solution $U(x,\mu)\in{\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ which is of at most quadratic growth, and the following SDEs admit unique solutions in $L^{2}_{r}$ for each $x\in\mathbb{R}$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu)$:
@@
Since $U$ is of at most quadratic growth, taking the limit as $T\rightarrow+\infty$ yields